Use the discriminant to determine whether the following equations have solutions that are two different rational solutions two different irrational solutios exactly one rational solution or two different imaginary solutions. 25xsquared=10x-1
Answer by mananth(16949)

25x^2-10+1=0
a=25, b=-10,c=1
b^2-4ac=100-100
b^2-4ac=0
One real solution